(a) What percentage of arginine side chains are protonated at a pH value of 11.5? (b) What is the average charge of the side chain of arginine at this pH? (pKa = 12)
Answer #1

a)

pH= pKa + log [A-] /[HA]

11.5-12 = log [A-] /[HA]

[A-] /[HA] = 0.316

[HA]/ [A-] = 3.16; That means 31.6 percent is protonated.

b)

When pH≠pKa we determine the mole fraction, for each charged species.

χCOO−=1/ (1+10pKa−pH) = 0.240269, this charge is negative ; remember this

χNH3+=1/(1+10pH−pKa) = 0.7597469

Net charge is their sum = 0.519477

Plugging in the numbers, we get:

χCOO−≈0.999981χCOO−≈0.999981

χNH+3≈0.993872